OpenStudy (anonymous):

A circle has the equation x^2+y^2+x-6y+9=0. Find the center and any x or y intercepts. I know it's a complete the square problem, but I'm messing up somewhere :/

OpenStudy (anonymous):

We know: \[ x^2+y^2+x-6y+9=0 \]So, we complete the square as such: \[ x^2+y^2+x-6y+9=x^2+x+y^2-6y+9=\\ \left(x+\frac{1}{2}\right)^2+(y-3)^2=\left(\frac{1}{2}\right)^2=\frac{1}{4} \]

OpenStudy (anonymous):

I think you should be able to work from there.

OpenStudy (anonymous):

Okay that's actually close to what I had, let's work backwards to see where I messed up. Thanks again! Hahah you've helped me multiple times on this, sorry I'm slow :)

OpenStudy (anonymous):

Sure thing

OpenStudy (anonymous):

Got it!

OpenStudy (anonymous):

Nice